OpenStudy (annie12m):

Medal to check my work: 5x + 10 = −4x − 17 +17 +17 =27 5x + 27 = -4x /5 /5 =-0.8x 27 = -0.8x /-0.8 /-0.8 = -33.75 -33.75 = x Correct? Or what I do wrong?

OpenStudy (annie12m):

@JeruDavis @laila22 @Mehek14 @neonumbrella5115

OpenStudy (neonumbrella5115):

So you see where you divided by 5? It would be much easier to subtract 5x from both sides

pooja195 (pooja195):

It's incorrect, 5x + 27 = -4x Start again from here don't divide

OpenStudy (annie12m):

5x + 27 = -4x -5 -5 = -9x 27 = -9x +9 +9 36 = x?

pooja195 (pooja195):

You would add 9 since you are wanting to isolate x you would divide both sides by 9 27 = -9x

OpenStudy (annie12m):

27 = -9x /-9 /-9 -3 = x?

pooja195 (pooja195):

perfect! :)

OpenStudy (annie12m):

thank you!

pooja195 (pooja195):

yw ^_^